In a study among kindergarten pupils of s certain private school, the average IQ is 100. A researcher studied the effect of his manipulative device on the intelligence of the children. A sample of 25 students who used the manipulatives has a mean of 140 with a standard deviation of 20. If the IQ of the students follows a normal distribution, what is the t-value?


1
Expert's answer
2022-05-03T11:33:02-0400
"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{140-100}{20/\sqrt{25}}=10"